OpenStudy (anonymous):

Suppose you are climbing a hill whose shape is given by the equation z = 1000 − 10x^2 − 5y^2 , where x, y, and z are measured in metres. You are standing at the point x = 5, y = 10, z = 250. (a) In what direction is the slope largest? (b) What is the slope in that direction? (c) In which directions is the slope 0?

OpenStudy (roxy_martinez):

Which equation can be rewritten as a linear equation after combining all like terms? a) x^2+8x+5=(x^2+4) b) x+4/x=9x+4 c) 2x^3+4x+5(x^3+4)-7x^3+4(x^3+100) d) 4^2+4x+5=5x^2+4 help in this question